# Induced three-neutron interactions with low cutoffs for dilute neutron matter

### v1

@@ Abstract
The properties of dilute neutron matter are mostly determined by the $s$-wave two-body (2N) interaction, while three-body (3N) interactions are suppressed by the Pauli principle. In a previous work, we showed that it can be advantageous to use the renormalization group based effective interactions with cutoffs scaled with the Fermi momentum, especially at low densities. In that case, induced 3N interactions may become important. In this work, we compute the 3N interaction induced by the similarity renormalization group flow of the $s$-wave 2N interaction. We work in the momentum-space hyperspherical partial wave basis and investigate its convergence properties. Then we study the effect of the induced 3N interaction on the equation of state of dilute neutron matter. We observe that the cutoff dependence of the equation of state is strongly reduced when the effect of induced 3N interaction is included.
@@ I Introduction
A microscopic theory of the inner crust of neutron stars is crucial for understanding their thermal evolution. Neutron matter in the inner crust is in the superfluid state, which is a necessary ingredient to explain the phenomenon of the observed glitches Chamel and Haensel (2008). In addition, the neutron gas is strongly correlated, although dilute (at the densities relevant to the inner crust). Our objective is to obtain reliable dilute neutron matter properties, such as the equation of state (EoS) and the pairing gap from microscopic calculations.
@@
An alternative is to use the so-called softened interactions such as the renormalization group based low-momentum interactions ($V_{\text{low-}k}$) or the similarity renormalization group (SRG) interactions Bogner et al. (2003, 2007a, 2010). In addition to decoupling the low- and high-momentum regions and softening the interactions, these have an important additional property: as one decreases the renormalization scale (cutoff) $\Lambda$ towards zero, the 2N matrix elements flow to the scattering length $a$.
@@
It has been demonstrated in Ramanan and Urban (2018) that the correct low-density limit for the superfluid critical temperature is obtained with a density-dependent cutoff $\Lambda=fk_{F}$, where $k_{F}$ is the Fermi momentum and $f$ is a scale factor, due to the effective resummation of the ladders, inherent in the flow towards the scattering length as $\Lambda\to 0$. In our previous work Palaniappan et al. (2023), we have shown that it is possible to obtain the correct EoS of neutron matter at very low densities using a 2N $V_{\text{low-}k}$ interaction with a density-dependent cutoff. In addition, the Bogoliubov many-body perturbation theory (BMBPT) corrections to the Hartree-Fock-Bogoliubov (HFB) ground state energy obtained using a density-dependent cutoff have better convergence compared to those obtained using a fixed cutoff. However, the HFB+BMBPT results in Palaniappan et al. (2023) show a small residual cutoff dependence, which indicates missing corrections.
@@
It is known that as one decreases the cutoff, three- and higher-body interactions are induced. The SRG presents an efficient method to consistently evolve two- and higher-body interactions. One possible choice for the 3N momentum space basis is the Jacobi partial-wave basis which has been used in the SRG-evolved 3N interactions in the study of both finite nuclei as well as infinite matter Hebeler (2012); Hebeler and Furnstahl (2013); Hebeler (2021). An alternative momentum space basis for the 3N interaction is the hyperspherical harmonic or hyperspherical partial-wave (HPW) basis, as was first demonstrated and applied to the triton in Wendt (2013a).
@@
In the present paper, we focus on the effect of the induced 3N interactions with low cutoffs on dilute neutron matter EoS, using the SRG formalism. We restrict ourselves to the 3N interactions induced by the 2N $s$-wave interaction, since this is by far the dominant channel at low densities. In Section II, we obtain the induced 3N interaction in the momentum space HPW basis using the SRG formalism. In Sec. III, we discuss our results for the induced 3N matrix elements, their convergence in the HPW basis, and their effect on the EoS of dilute neutron matter, in particular the cutoff dependence. In Sec. IV, we summarize our current work and present an outlook for possible future studies. Technical details are provided in the appendices.
@@ II.1 Similarity Renormalization Group
The SRG flow equation is given by
@@
where $\eta_{s}$ is the generator of the unitary transformation, $H_{s}$ is the Hamiltonian, and $s$ is the flow parameter. The flow parameter $s$ is usually replaced with the decoupling scale $\lambda\equiv s^{-1/4}$. It is well known that the flow equation induces higher-body interactions as one evolves in $s$.
Since we are interested in dilute neutron matter, we concentrate on the 3N interaction induced by the 2N $s$-wave (${}^{1}S_{0}$) interaction. The “bare” (or “genuine”) 3N interaction and 3N interactions induced by higher partial waves, as well as induced 4N (and higher) interactions, will be neglected. In an induced 3N interaction, one of the three particles interacts consecutively with the other two particles via 2N interactions. If the 2N interactions are restricted to the $s$-wave channel, then this particle must have an opposite spin with respect to the other two particles. And since the $s$-wave interaction does not flip the spins, the incoming and outgoing particles for the induced 3N interaction must have the same spins. Fig. 1 shows a schematic illustration of the 2N and 3N terms in the Hamiltonian.
@@
In the latter equation, we used the following definitions and notations: $A^{(ij)}=I-P^{(ij)}$ is proportional to the antisymmetrizer with respect to particles $i$ and $j$. To be specific, $I_{1231^{\prime}2^{\prime}3^{\prime}}=\delta_{11^{\prime}}\delta_{22^{\prime}}\delta_{33^{\prime}}$ is the 3N identity operator and $P^{(13)}_{1231^{\prime}2^{\prime}3^{\prime}}=\delta_{13^{\prime}}\delta_{22^{\prime}}\delta_{31^{\prime}}$ is the permutation operator of particles 1 and 3. Notations of the form $(AB)_{1231^{\prime}2^{\prime}3^{\prime}}$ should be understood as operator products $\sum_{1^{\prime\prime}2^{\prime\prime}3^{\prime\prime}}A_{1231^{\prime\prime}2^{\prime\prime}3^{\prime\prime}}B_{1^{\prime\prime}2^{\prime\prime}3^{\prime\prime}1^{\prime}2^{\prime}3^{\prime}}$. $V^{(ij)}$ denotes the 2N interaction between particles $i$ and $j$ “embedded” in the 3N space. For example,
@@ II.2 Two-body flow equation
$\frac{d}{ds}V_{s}(k,k^{\prime})=$ $-(k^{2}-k^{\prime 2})^{2}V_{s}(k,k^{\prime})$
@@
×$\times V_{s}(k,k^{\prime\prime})V_{s}(k^{\prime\prime},k^{\prime}).$ (7)
@@ II.3 Three-body flow equation
×$\hskip 14.22636pt\times V_{0}\bigg{(}K\sqrt{\frac{1+x}{2}},K^{\prime}\sqrt{\frac{1+x^{\prime}}{2}}\bigg{)},$ (11)
where $x^{\prime}=1-\big{(}\frac{K}{K^{\prime}}\big{)}^{2}(1-x)$ Wendt (2013b). In this equation, $N^{l_{2}l_{1}}_{L}$ is a normalization factor (see Appendix C) and $P^{(\alpha,\beta)}_{n}$ denotes a Jacobi polynomial. For $K<K^{\prime}$, the integral over $x$ runs from $-1$ to $1$ and can be computed using Gauss-Jacobi quadrature. For $K>K^{\prime}$, one can use $V_{lLL^{\prime}}(K,K^{\prime})=V_{lL^{\prime}L}(K^{\prime},K)$.
@@
Unlike the last three ($VW$, $WV$, and $WW$) terms, we have not written the $VV$ term as a sum over $L^{\prime\prime}$. The reason for this is that in the $VV$ term, the $L^{\prime\prime}$ summation converges very slowly at low momenta. In fact, combining these spectator delta functions with the momentum-conserving delta functions of the 2N matrix elements $V$, one can see that in the $VV$ term of Eq. (4), the internal momenta $789$ can be expressed in terms of the external momenta $123456$, and there is actually no summation over intermediate states at all. Then, the $VV$ term can be computed from the following expression (refer to Appendix D for more details on the derivation):
@@ III Results and Discussion
Since we truncate the intermediate $L^{\prime\prime}$ sum in Eq. (II.3) at some finite value $L_{\text{max}}$, it is necessary to check that the matrix elements are converged. Fig. 4 shows the convergence of the $W_{111}$ matrix elements. Convergence is reached when channels with $L^{\prime\prime}$ up to $L_{\text{max}}=13$ are included in the summation. We emphasize that this is only possible because we use Eq. (II.3) for the $VV$ term. If instead of using Eq. (II.3), the $VV$ term is computed using the embedded matrix elements along with the intermediate $L^{\prime\prime}$ summation, the convergence of the matrix elements at low momenta is much slower.
@@
Since we use a density-dependent cutoff to get the correct low-density limit of the neutron matter EoS Palaniappan et al. (2023), we have to flow the matrix elements to much lower cutoffs (up to $\lambda=0.1\,\text{fm}{}^{-1}$) than it is usually done. When the cutoff gets very small compared to the last point in the momentum mesh $k_{\text{max}}$ (say, $\lambda=k_{\text{max}}/5$), we change the momentum mesh in the following manner: (1) make the mesh denser at lower momenta ($k,k^{\prime}\lesssim\lambda$), (2) cut the upper edge of the mesh ($k,k^{\prime}>3.8\lambda$), where there is only a narrow tail of non-zero matrix elements along the diagonal, and (3) apply a smooth regulator at the upper edge of the new mesh to avoid numerical instabilities (due to step (2)) as we flow further down. Since the tail being excluded is almost diagonal, step (2) has no effect on the results Bogner et al. (2007b).
@@ III.2 Induced 3N contribution to the neutron-matter EoS
$\varepsilon_{\text{3HF}}=\frac{\braket{\text{HF}}{W}{\text{HF}}}{\mathcal{V}}=\frac{1}{2\mathcal{V}}\sum_{p_{1},p_{2},p_{3}<k_{F}}W_{123123}\,.$ (15)
The integrals can be computed numerically using Monte-Carlo integration. The dashed line in Figure 5 shows the result in units of the ideal Fermi gas energy [^1] The symbols $\varepsilon$ and $E$ denote, respectively, the energy density and the energy per particle, related by $E=\varepsilon/n$, where $n=k_{F}^{3}/(3\pi^{2})$ is the density. [$\varepsilon_{\text{FG}}=k_{F}^{5}/(10\pi^{2})$] as a function of the Fermi momentum, for the choice $\lambda=2k_{F}$. As we can see, it is small and repulsive. For a bare 3N interaction, one would expect it to contribute only at high density but for the induced interaction, it is completely different. The reason is the density-dependent cutoff: the smaller the $k_{F}$, the smaller $\lambda$ gets, and the induced 3N matrix elements get bigger.
However, in our previous work Palaniappan et al. (2023), we have seen that, especially at very low density, the HF ground state is not the best starting point. There, the starting point was the HFB ground state to which perturbative corrections were added in BMBPT. Therefore, let us see how the contribution of the induced 3N interaction changes if we compute it in first-order perturbation theory by taking the expectation value in the HFB ground state instead of the HF one. In this exploratory study, we do not include the effect of the induced 3N interaction on the HFB ground state or its BMBPT corrections (similar to Fig. 49 of Ref. Hebeler (2021)). Note that the 2N calculations presented in this and the next subsection are obtained with the full SRG evolved 2N interaction (including partial waves with $j\leq 6$ yields converged results for the density range $k_{F}\leq 1.4\,\text{fm}{}^{-1}$), while for the 3N calculations, we consider only the contribution of the $s$-wave-induced 3N interaction, as explained before.
@@
In the limit of no pairing, the second term vanishes, the $v_{i}$ become step functions $\theta(k_{F}-p_{i})$, and the first term reduces to the expression of the expectation value in the HF ground state, Eq. (15). The result is shown in Fig. 5 as the solid line.
Comparing the HF and HFB expectation values, we see that they are identical at $k_{F}\gtrsim 0.9\,\text{fm}{}^{-1}$ when the gap becomes small compared to the Fermi energy (similar to the HF and HFB energies in the 2N case, cf. Fig. 4 in Ref. Palaniappan et al. (2023)). While the first term in the right-hand side of Eq. (III.2) makes a similar contribution as in the HF case, the second term gives a significant contribution at low densities ($k_{F}\lesssim 0.7\,\text{fm}{}^{-1}$). The physical interpretation of this second term is the increase in energy due to the reduction of the pairing gap if one includes an additional contribution to the pairing interaction from the 3N interaction by closing one line.
Coming back to the question of convergence, Fig. 6(a) shows the convergence of the 3N contribution with respect to $L$ with a fixed $l=1$ and Fig. 6(b) shows the convergence of the same with respect to $l$. The energy has converged with the inclusion of hyperspherical partial waves up to $L_{\text{max}}=13$ and $l_{\text{max}}=5$.
@@
Figure 7: Cutoff Dependence of the EoS of dilute neutron matter with the 2N (BMBPT3) interactions (short-dashed lines) only, with the 2N (BMBPT3) + induced 3N (HF) interactions (long-dashed lines) and with the 2N (BMBPT3) + induced 3N (HFB) interactions (solid lines): (a) The ground state energy as a function of the Fermi momenta for different values of the scale factor $f=\lambda/k_{F}$, (b) The ground state energy as a function of the scale factor $f=\lambda/k_{F}$ for different values of the Fermi momenta. All energies are shown in units of the energy of the ideal Fermi gas.

### v2

@@ Abstract
The properties of dilute neutron matter are mostly determined by the $s$-wave two-body (2N) interaction, while three-body (3N) interactions are suppressed by the Pauli principle. In a previous work, we showed that it can be advantageous to use renormalization group based effective interactions with cutoffs scaled with the Fermi momentum, especially at low densities. In that case, induced 3N interactions may become important. In this work, we compute the 3N interaction induced by the similarity renormalization group flow of the $s$-wave 2N interaction. We work in the momentum-space hyperspherical partial wave basis and investigate its convergence properties. Then we study the effect of the induced 3N interaction on the equation of state of dilute neutron matter. We observe that the cutoff dependence of the equation of state is strongly reduced when the effect of induced 3N interaction is included.
@@ I Introduction
The superfluid unbound neutrons in the inner crust of neutron stars are crucial for understanding their thermal evolution and the phenomenon of the observed pulsar glitches Chamel and Haensel (2008). In addition, the neutron gas is strongly correlated, although dilute. Our objective is to obtain reliable dilute neutron matter properties, such as the equation of state (EoS) and the pairing gap from microscopic calculations. This would allow us to better constrain the energy density functionals, which are used in calculations of the inner crust, but do not have the correct behavior at relevant low densities Yang et al. (2016); Gupta and Urban (2024); Grams and Margueron (2024), and pairing-related quantities such as the superfluid density, the heat capacity, and neutrino emission rates Sedrakian and Clark (2019).
@@
An alternative is to use the so-called softened interactions such as the renormalization group based low-momentum interactions ($V_{\text{low-}k}$) or the similarity renormalization group (SRG) interactions Bogner et al. (2003, 2007a, 2010). In addition to decoupling the low- and high-momentum regions and softening the interactions, these have an important additional property: as one decreases the renormalization scale (cutoff) $\Lambda$ towards zero, the 2N matrix element $V_{{}^{1}S_{0}}(0,0)$ flows to the scattering length $a$. Taking the cutoff $\Lambda\sim 2k_{\text{F}}$ makes the effective interaction perturbative as discussed in Bogner et al. (2005).
@@
It has been demonstrated in Ramanan and Urban (2018) that the correct low-density limit for the superfluid critical temperature is obtained with a density-dependent cutoff $\Lambda=fk_{\text{F}}$, where $k_{\text{F}}$ is the Fermi momentum and $f$ is a scale factor, due to the effective resummation of the ladders, inherent in the flow towards the scattering length as $\Lambda\to 0$ Bogner et al. (2005). In our previous work Palaniappan et al. (2023), we have shown that it is possible to obtain the correct EoS of neutron matter at very low densities using a 2N $V_{\text{low-}k}$ interaction with a density-dependent cutoff and a scale factor $f\sim 1.5-2.5$. In addition, the Bogoliubov many-body perturbation theory (BMBPT) corrections to the Hartree-Fock-Bogoliubov (HFB) ground state energy obtained using a density-dependent cutoff have better convergence compared to those obtained using a fixed cutoff. However, the HFB+BMBPT results in Palaniappan et al. (2023) show a small residual cutoff dependence, which indicates missing corrections. It is known that as one decreases the cutoff, three- and higher-body interactions are induced Bogner et al. (2010). However, these were not included in our previous work, since it is not straight-forward to extend $V_{\text{low-}k}$ to higher-body interactions.
The SRG presents an efficient method to consistently evolve two- and higher-body interactions. Because of the Pauli principle, the 3N contact term, which should be dominant in symmetric matter at very low density, is absent in pure neutron matter Hebeler and Schwenk (2010). The leading “bare” (or “genuine”) 3N force for three neutrons is the two-pion exchange contribution which starts to contribute only at higher densities. Therefore, in the present paper, we neglect the bare 3N force and consider only the one induced by the SRG flow. Furthermore, we restrict ourselves to the 3N interactions induced by the 2N $s$-wave interaction, since this is by far the dominant channel at low densities.
@@
Concerning the practical implementation, one possible choice for the 3N momentum space basis is the Jacobi partial-wave basis which has been used in the SRG-evolved 3N interactions in the study of both finite nuclei as well as infinite matter Hebeler (2012); Hebeler and Furnstahl (2013); Hebeler (2021). An alternative momentum space basis for the 3N interaction is the hyperspherical harmonic or hyperspherical partial-wave (HPW) basis, as was first demonstrated and applied to the triton in Wendt (2013a).
@@
In Section II, we obtain the induced 3N interaction in the momentum space HPW basis using the SRG formalism. In Sec. III, we discuss our results for the induced 3N matrix elements, their convergence in the HPW basis, and their effect on the EoS of dilute neutron matter, in particular the cutoff dependence. In Sec. IV, we summarize our current work and present an outlook for possible future studies. Technical details are provided in the appendices.
@@ II.1 Similarity Renormalization Group
The SRG flow equation is given by Bogner et al. (2007a)
@@
where $\eta_{s}$ is the generator of the unitary transformation, $H_{s}$ is the Hamiltonian, and $s$ is the flow parameter. The flow parameter $s$ is usually replaced with the decoupling scale $\lambda\equiv s^{-1/4}$. It is well known that the flow equation induces higher-body interactions as one evolves in $s$ Bogner et al. (2007a, 2010).
Since we are interested in dilute neutron matter, we concentrate on the 3N interaction induced by the 2N $s$-wave (${}^{1}S_{0}$) interaction and neglect 3N interactions induced by higher partial waves. Further, the bare 3N interaction as well as induced 4N (and higher) interactions are also neglected. In an induced 3N interaction, one of the three particles interacts consecutively with the other two particles via 2N interactions. If the 2N interactions are restricted to the $s$-wave channel, then this particle must have an opposite spin with respect to the other two particles. And since the $s$-wave interaction does not flip the spins, the incoming and outgoing particles for the induced 3N interaction must have the same spins. Fig. 1 shows a schematic illustration of the 2N and 3N terms in the Hamiltonian.
@@
where $\varepsilon_{ij}=\varepsilon_{i}+\varepsilon_{j}$, and $\varepsilon_{ijk}=\varepsilon_{i}+\varepsilon_{j}+\varepsilon_{k}$. In Eq. (4), we used the following definitions and notations: $A^{(ij)}=I-P^{(ij)}$ is proportional to the antisymmetrizer with respect to particles $i$ and $j$. To be specific, $I_{1231^{\prime}2^{\prime}3^{\prime}}=\delta_{11^{\prime}}\delta_{22^{\prime}}\delta_{33^{\prime}}$ is the 3N identity operator and $P^{(13)}_{1231^{\prime}2^{\prime}3^{\prime}}=\delta_{13^{\prime}}\delta_{22^{\prime}}\delta_{31^{\prime}}$ is the permutation operator of particles 1 and 3. Notations of the form $(AB)_{1231^{\prime}2^{\prime}3^{\prime}}$ should be understood as operator products $\sum_{1^{\prime\prime}2^{\prime\prime}3^{\prime\prime}}A_{1231^{\prime\prime}2^{\prime\prime}3^{\prime\prime}}B_{1^{\prime\prime}2^{\prime\prime}3^{\prime\prime}1^{\prime}2^{\prime}3^{\prime}}$. $V^{(ij)}$ denotes the 2N interaction between particles $i$ and $j$ “embedded” in the 3N space. For example,
@@ II.2 Two-body flow equation
$\frac{d}{ds}{V}(k,k^{\prime})=$ $-(k^{2}-k^{\prime 2})^{2}{V}(k,k^{\prime})$
@@
×$\times{V}(k,k^{\prime\prime}){V}(k^{\prime\prime},k^{\prime}).$ (7)
@@ II.3 Three-body flow equation
×$\hskip 14.22636pt\times{V}\bigg{(}K\sqrt{\frac{1+x}{2}},K^{\prime}\sqrt{\frac{1+x^{\prime}}{2}}\bigg{)},$ (11)
where $x^{\prime}=1-\big{(}\frac{K}{K^{\prime}}\big{)}^{2}(1-x)$ Wendt (2013b). In this equation, $N^{l_{2}l_{1}}_{L}$ is a normalization factor (see Appendix C) and $P^{(\alpha,\beta)}_{n}$ denotes a Jacobi polynomial. For $K\leq K^{\prime}$, the integral over $x$ runs from $-1$ to $1$ and can be computed using Gauss-Jacobi quadrature. For $K>K^{\prime}$, one can use $V_{lLL^{\prime}}(K,K^{\prime})=V_{lL^{\prime}L}(K^{\prime},K)$.
@@
Unlike the last three ($VW$, $WV$, and $WW$) terms, we have not written the $VV$ term as a sum over $L^{\prime\prime}$. The reason for this is that in the $VV$ term, the $L^{\prime\prime}$ summation converges very slowly at low momenta. In fact, combining the spectator delta functions with the momentum-conserving delta functions of the 2N matrix elements $V$, one can see that in the $VV$ term of Eq. (4), the internal momenta $789$ can be expressed in terms of the external momenta $123456$, and there is actually no summation over intermediate states at all. Then, the $VV$ term can be computed from the following expression (refer to Appendix D for more details on the derivation):
@@ III Results and Discussion
Since we truncate the intermediate $L^{\prime\prime}$ sum in Eq. (II.3) at some finite value $L_{\text{max}}$, it is necessary to check that the matrix elements have converged. Fig. 4 shows the convergence of the $W_{111}$ matrix elements. Convergence is reached when channels with $L^{\prime\prime}$ up to $L_{\text{max}}=13$ are included in the summation. We emphasize that this is only possible because we use Eq. (II.3) for the $VV$ term. If instead of using Eq. (II.3), the $VV$ term is computed using the embedded matrix elements along with the intermediate $L^{\prime\prime}$ summation, the convergence of the matrix elements at low momenta is much slower.
@@
Since we use a density-dependent cutoff to obtain the correct low-density limit of the neutron matter EoS Palaniappan et al. (2023), we have to flow the matrix elements to much lower cutoffs than is usually done. Figure 5 shows the diagonal ${}^{1}S_{0}$ matrix elements for different cutoffs. We note that they flow to $a$ (for $k\ll 1/a$) as one lowers the cutoff, similar to the ${}^{1}S_{0}$ $V_{\text{low-}k}(0,0)$ matrix elements as shown in Fig. 15 in Bogner et al. (2010). [^1] This limit is in practice reached only at $\lambda\ll 1/a$ since $1/V_{{}^{1}S_{0}}(0,0)\approx 1/a-2\lambda/\pi$ Ramanan and Urban (2018); Urban and Ramanan (2021) as one can easily understand in the $V_{\text{low-}k}$ case.
Figure 5: Diagonal ${}^{1}S_{0}$ matrix elements for different SRG cutoffs. The dashed and solid lines correspond to matrix elements computed with and without readjusting the mesh, as explained in the text. The gray long dashed line corresponds to the ${}^{1}S_{0}$ scattering length $a=-18.95\,\text{fm}$ Entem et al. (2017).
While using a fine and large momentum mesh as in Fig. 5 is feasible for demonstration purposes, it would be too time and memory consuming to do this for all partial waves, especially in the 3N case. Hence, we restrict the momentum mesh to the relevant range by readjusting it during the evolution. When the cutoff gets very small compared to the last point in the momentum mesh $k_{\text{max}}$ (say, $\lambda=k_{\text{max}}/5$), we change the momentum mesh in the following manner: (1) make the mesh denser at lower momenta ($k,k^{\prime}\lesssim\lambda$), (2) cut the upper edge of the mesh ($k,k^{\prime}>3.8\lambda$), where there is only a narrow tail of non-zero matrix elements along the diagonal, and (3) apply a smooth regulator at the upper edge of the new mesh to avoid numerical instabilities (due to step (2)) as we flow further down. Since the tail being excluded is almost diagonal, step (2) has no effect on the results Bogner et al. (2007b). This can be seen in Fig. 5 from the perfect agreement of the matrix elements for $k\lesssim 3.8\lambda$ computed with (dashed lines) and without (solid lines) readjusting the mesh.
@@ III.2 Induced 3N contribution to the neutron-matter EoS
$\varepsilon_{\text{3HF}}=\frac{\braket{\text{HF}}{W}{\text{HF}}}{\mathcal{V}}=\frac{1}{2\mathcal{V}}\sum_{p_{1},p_{2},p_{3}<k_{\text{F}}}W_{123123}\,.$ (15)
The integrals can be computed numerically using Monte-Carlo integration. The dashed line in Figure 6 shows the result in units of the ideal Fermi gas energy [^2] The symbols $\varepsilon$ and $E$ denote, respectively, the energy density and the energy per particle, related by $E=\varepsilon/n$, where $n=k_{\text{F}}^{3}/(3\pi^{2})$ is the density. [$\varepsilon_{\text{FG}}=k_{\text{F}}^{5}/(10\pi^{2})$] as a function of the Fermi momentum, for the choice $\lambda=2k_{\text{F}}$. As we can see, it is small and repulsive. For a bare 3N interaction, one would expect it to contribute only at high density, but for the induced interaction, it is completely different. The reason is the density-dependent cutoff: the smaller the $k_{\text{F}}$, the smaller $\lambda$ gets, and the induced 3N matrix elements get bigger. However, at very small $k_{\text{F}}$ values, the ratio $\varepsilon_{3\text{HF}}/\varepsilon_{\text{FG}}\propto k_{\text{F}}^{4}\braket{W}$ starts to decrease because the matrix elements at momenta of order $k_{\text{F}}$ grow slower than $1/\lambda^{4}$.
However, in our previous work Palaniappan et al. (2023), we have seen that, especially at very low density, the HF ground state is not the best starting point. There, the starting point was the HFB ground state to which perturbative corrections were added in BMBPT. Therefore, let us see how the contribution of the induced 3N interaction changes if we compute it in first-order perturbation theory by taking the expectation value in the HFB ground state instead of the HF one. In this exploratory study, we do not include the effect of the induced 3N interaction on the HFB ground state or its BMBPT corrections (similar to Fig. 49 of Ref. Hebeler (2021)). Note that the 2N calculations presented in this and the next subsection are obtained with the full SRG evolved 2N interaction (including partial waves with $j\leq 6$ yields converged results for the density range $k_{\text{F}}\leq 1.4\,\text{fm}{}^{-1}$), while for the 3N calculations, we consider only the contribution of the $s$-wave-induced 3N interaction, as explained before.
@@
In the limit of no pairing, the second term vanishes, the $v_{i}$ become step functions $\theta(k_{\text{F}}-p_{i})$, and the first term reduces to the expression of the expectation value in the HF ground state, Eq. (15). The result is shown in Fig. 6 as the solid line.
Comparing the HF and HFB expectation values, we see that they are identical at $k_{\text{F}}\gtrsim 0.9\,\text{fm}{}^{-1}$ when the gap becomes small compared to the Fermi energy (similar to the HF and HFB energies in the 2N case, cf. Fig. 4 in Ref. Palaniappan et al. (2023)). While the first term in the right-hand side of Eq. (III.2) makes a similar contribution as in the HF case, the second term gives a significant contribution at low densities ($k_{\text{F}}\lesssim 0.7\,\text{fm}{}^{-1}$). The physical interpretation of this second term is the increase in energy due to the reduction of the pairing gap if one includes an additional contribution to the pairing interaction from the 3N interaction by closing one line.
Coming back to the question of convergence, Fig. 7(a) shows the convergence of the 3N contribution with respect to $L$ with a fixed $l=1$ and Fig. 7(b) shows the convergence of the same with respect to $l$. The energy has converged with the inclusion of hyperspherical partial waves up to $L_{\text{max}}=13$ and $l_{\text{max}}=5$. It is interesting to note that $l=1$ is the dominant contribution, and not $l=0$. This is because $l=1$ contains $L=L^{\prime}=1$, whereas $l=0$ starts only from $L=L^{\prime}=2$ due to antisymmetrization, which forbids $L$ or $L^{\prime}=0$. Also, note that odd and even $l$ contributions have opposite signs.
@@
Figure 8: Cutoff Dependence of the EoS of dilute neutron matter with the 2N (BMBPT3) interactions (short-dashed lines) only, with the 2N (BMBPT3) + induced 3N (HF) interactions (long-dashed lines) and with the 2N (BMBPT3) + induced 3N (HFB) interactions (solid lines): (a) The ground state energy as a function of the Fermi momenta for different values of the scale factor $f=\lambda/k_{\text{F}}$, (b) The ground state energy as a function of the scale factor $f=\lambda/k_{\text{F}}$ for different values of the Fermi momenta. All energies are shown in units of the energy of the ideal Fermi gas.